# Evaluating LLM-corrupted Crowdsourcing Data Without Ground Truth

### v1

@@ 5.3 Assumption Verification
The first two hypotheses correspond to 4.1 and 4.2 respectively and the other two hypotheses correspond to 4.5. More broadly, our results in this section provide valuable insights into the correlations between different LLMs and their correlations compared with human labelers in the context of subjective labeling questions.
@@
For each considered LLM, we prompt it to independently generate responses to each question in the dataset three times using the default temperature. These responses are treated as samples of the cheap signals $Z_{i}$, $Z_{j}$, and $Z$. Furthermore, we randomly select two human responses in the original dataset as samples of $X_{i}$ and $X_{j}$, respectively. Using these samples across all tasks, we estimate the conditioned joint distributions and then compute the conditioned mutual information. We report the results for $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$ on two datasets in Fig. 2, while results for other datasets and for $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$ are deferred to Section D.1.
(a) Hatefulness
(b) Preference Alignment
@@
Figure 2: The TVD mutual information between $Z_{i}$ and $X_{j}$ conditioned on $Z$, i.e. $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$. LLM names on the x-axis denote the models used to sample $Z$ while each bar in the same group represents the model used to sample $Z_{i}$. We use “Random Cheaters” as a reference, who randomly selects an answer to each question according to the prior.
@@
Hypotheses 1 and 2 generally do not hold.  In Fig. 2, we observe that when $Z_{i}$ and $Z$ are sampled from different models, the mutual information between $Z_{i}$ and $X_{j}$ is often significantly higher than that expected from the random reporting (red bar). Even when $Z_{i}$ and $Z$ are independently sampled by the same model, conditioning on $Z$ reduces but does not eliminate this correlation. This indicates that independently generated responses from the same LLM can remain dependent after conditioning on another independent sample, contradicting Hypothesis 1. As presented in Section D.1, this pattern continues to hold for $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$, rejecting hypothesis 2.
@@
Hypothesis 3 and 4 generally hold when conditioning on the “right” model.  Encouragingly, the conditioned correlation between an LLM and a human labeler is consistently weaker than that between two human labelers when $Z_{i}$ and $Z$ are sampled from the same model. For example, in Fig. 2(b), the unconditioned mutual information between GPT-4-generated labels and human labels initially exceeds that of two expert labels. However, after conditioning on independent GPT-4 responses, $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$ becomes significantly higher than $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$. This pattern is consistent across all tested datasets, providing strong support for hypothesis 3. Results for Hypothesis 4 are deferred to the appendix.
@@
Our empirical results suggest that as of the time of this study and based on the subjective labeling tasks we tested, human labelers exhibit distinct correlation patterns that no LLM replicates.
5.4 Detecting Cheating Agents
@@
In this subsection, we focus on identifying cheating agents with naive strategies and comparing various methods based on their area under the ROC curve (AUC).
@@ 5.4.1 Simulating Noisy Crowds
Real-world crowdsourcing datasets do not include labels identifying whether an agent is cheating or working diligently. To enable evaluation, we simulate cheating agents and treat the original dataset labels as representing “high-effort” responses. Specifically, we replace a randomly selected fraction of labelers in the original dataset with cheaters adopting naive strategies. To reflect the noisy nature of real-world crowdsourcing data, we simulate three types of cheaters, each representing a common naive strategy observed in practice (Yang et al., 2019; Becker et al., 2021):
@@
    An $\alpha_{llm}$ fraction of labelers are LLM cheaters, who report the LLM-generated labels on all assigned questions.
@@
    An $\alpha_{r}$ fraction of labelers are random cheaters, who generate labels by independently sampling from the marginal distribution over labels for each assigned question.
@@
    An $\alpha_{b}$ fraction of labelers are biased cheaters, who report the dataset’s majority label on 90% of questions and choose uniformly at random on the rest.
Figure 3: Distribution of the conditioned CA scores with $(\alpha_{llm},\alpha_{r},\alpha_{b})=(0.1,0.05,0.05)$, where GPT-4 generates samples for $Z_{i}$, $Z_{j}$, and $Z$ using the preference alignment dataset.
Figure 3 shows the distribution of conditioned CA scores, where both the LLM cheaters’ responses and the principal’s signals are independently sampled from separate GPT-4 outputs. As shown, human workers achieve higher scores than any type of cheater on average. However, some workers fall below the average score of cheaters. Several factors contribute to this issue. For example, a subset of workers may themselves be cheaters in the dataset, or they may have been assigned many tasks that are inherently difficult to classify. Consequently, no detection method can achieve perfect accuracy. Next, we will examine how well our method distinguishes workers from cheaters compared to other baselines.
@@ 5.4.2 Baseline Measurements
We introduce the key ideas of the following four baselines.
@@
    We consider the Output Agreement (OA) mechanism that directly counts the frequency of agreements between two agent’s responses. Specifically, while scoring labeler $i$, we iteratively pair her with a different labeler $j$ and let $M_{i,j}$ denote the set of questions answered by both labelers. Then, the score of labeler $i$ is the average of the empirical probability of agreement, i.e. $\frac{1}{n}\sum_{j\in[n]\backslash\{i\}}\sum_{q\in M_{i,j}}\frac{\mathbbm{1}\left[\hat{X}_{i,q}=\hat{X}_{j,q}\right]}{|M_{i,j}|}$, and let $\frac{0}{0}=0$.
@@
    We use the Correlated Agreement (CA) mechanism as a baseline, which is introduced in Appendix A. The expected CA score is the unconditioned mutual information $I_{\text{TVD}}(\hat{X}_{i};\hat{X}_{j})$.
@@
    Another common method to estimate the reliability of the noisy agents is the Expectation-Maximization (EM) algorithm under the DS model (Dawid and Skene, 1979). Under the DS model, each agent $i$ makes mistakes according to a confusion matrix $\Gamma^{i}$ where $\Gamma^{i}_{h,l}$ denotes the probability that labeler $i$ reports $h$ when the ground truth answer is $l$. The EM algorithm first initializes a prior of the ground truth for each question and a confusion matrix for each agent. Then, it iteratively goes through an E-step and an M-step. In the E-step, the algorithm computes the posterior of the ground truth of each question given the prior and the confusion matrix. In the M-step, the confusion matrix of each agent is updated using the posterior computed in the E-step. The above two steps are iterated until the likelihood of the data matrix $\hat{X}$ converges. Finally, we score the reliability of agent $i$ as $\sum_{h\in\Sigma}\Gamma^{i}_{h,h}\tilde{P}(h)$ where $\tilde{P}(h)$ is the empirical marginal distribution of observing $h$ in the dataset.
@@
    Lastly, we consider the Conditioned Agreement ($\text{OA}_{Z}$) which counts the frequency of agreements only when the agreement is different from the principal’s sample on that question. Similar to the agreement metric, the score of labeler $i$ is the average of the empirical probability of agreement conditioned on $Z$, i.e. $\frac{1}{n}\sum_{j\in[n]\backslash\{i\}}\sum_{q\in M_{i,j}}\frac{\mathbbm{1}\left[\hat{X}_{i,q}=\hat{X}_{j,q}\&\hat{X}_{i,q}\neq Z_{q}\right]}{|M_{i,j}|}$. Intuitively, if labeler $i$ uses an LLM whose generated labels are similar to $Z$, then $\hat{X}_{i,q}\neq Z_{q}$ will happen with a small probability, meaning that the conditioned agreement score will be small.
We note that there are many more heuristic methods to estimate the reliability of crowdsourcing agents (Yin et al., 2008; Embretson and Reise, 2013; Piech et al., 2013). Many of these methods are variants of the EM algorithm under a different context-dependent model. We nonetheless only use the EM algorithm with the DS model to represent this line of methods because they all ignore the principal’s information $Z$ and thus are unlikely to perform well with LLM-corrupted datasets.
@@ 5.4.3 Results
(a) AUC v.s. fraction of LLM cheaters.
(b) AUC v.s. sampled questions.
Figure 4: The area under the ROC curve (AUC) for various methods on the preference alignment dataset as we vary (a) the fraction of LLM cheaters and (b) the proportion of questions the principal can sample. In (a), GPT-4 generates samples for $Z_{i}$, $Z_{j}$, and $Z$; in (b), AUC curves for all five LLM-cheater types are shown (omitting error bars for clarity).
Given a dataset $\hat{X}$ and samples of the cheap signal $Z$, we use our method as well as the baselines to score each labeler. We use the AUC score to measure the ability of each discussed metric to distinguish the simulated cheaters from the original human labelers.
Figure 4(a) presents a comparison of different measures while varying the fraction of LLM cheaters in the crowds. For each value of $\alpha_{llm}$ ranging from 00 to $0.2$, we randomly sample $\alpha_{r}$ and $\alpha_{b}$ uniformly from $[0,0.2]$ and report the average and error bars over $50$ such trials. Our result suggests that classic methods, such as the EM algorithm and the CA mechanism, perform well when the number of LLM cheaters is low. However, their AUCs degrade significantly as the fraction of LLM cheaters increases. On the other hand, the performance of $\text{OA}_{Z}$, which scores agents based on the similarity between their reports and $Z$, tends to have large variance and performs poorly when there are few LLM cheaters. In contrast, our proposed conditioned CA mechanism has the most robust performance for the mixed crowds and has the lowest variance.
Table 1 summarizes the results from Figure 4(a), reporting the average AUC score and the bottom 10% quantile for each method on each dataset w.r.t. the randomness of cheater fractions. A high average AUC reflects strong overall performance across random cheater compositions, while a high 10% quantile indicates robustness to the worst-case scenarios. The results support our earlier claim that our method consistently demonstrates the highest robustness across all tested settings.
Table 1: Average AUC (and bottom 10% quantiles) across datasets and methods with GPT-4 cheaters.
@@
Our experiments do not claim that our method outperforms all baselines in every scenario. Traditional methods like EM often perform better when all agents are random or biased cheaters. However, our method demonstrates the most robust performance across all settings due to its theoretical guarantees. This is particularly useful as, in practice, the designer usually has little clue about the composition of the crowds. In contrast, the baselines exhibit cases where the AUC drops below 0.5, meaning the algorithm completely misclassifies good workers as cheating agents.
Detection Performance vs. Number of Samples
We further examine how many questions the principal needs to sample to effectively detect cheaters. In Fig. 4(b), the AUC of the conditioned CA mechanism is shown with the LLM-cheater fraction fixed at 0.15, while the fraction of random and biased cheaters is each independently sampled from $[0,0.2]$. Similar plots for other datasets are provided in Section D.2.2. As the sampling rate increases from 20% to 100%, the average AUC also increases. While the results vary across the datasets and the LLMs that cheaters use, we observe that, except for the hatefulness labeling dataset with GPT-3.5 cheaters, sampling 50% of the questions is sufficient to outperform the baseline CA mechanism. Furthermore, in cases where human and AI labels are well separated, sampling just 20% can suffice.
@@
First, Section D.2.1 presents additional figures and tables for other datasets and LLMs, further supporting the claims made in the main body. Furthermore, when the principal is uncertain about which LLM the cheaters are using, we propose a method that computes the CA score conditioned on each candidate LLM’s outputs and uses the minimum score across them as the final score (Section D.2.3). Finally, in Appendix C and Section D.3, we present a heuristic extension of our method to open-ended textual responses. This serves as a preliminary step toward generalizing our approach beyond labeling tasks, with further refinement and large-scale validation left for future work.
6 Limitations
@@
We acknowledge several limitations of our method. First, our method is primarily designed for crowdsourcing datasets comprising subjective questions or tasks where human experts are considered the gold standard. If LLMs significantly surpass human performance, the correlations between human responses conditioned on $Z$ may no longer reflect the ground truth $Y$. In such cases, applying our method could mistakenly reward incorrect correlations.
@@
Second, as discussed in our experiment section, our method does not outperform all baselines in every scenario. In particular, when LLM responses are diverse, e.g. every cheating agent applies a different model or there are only a few cheating agents, heuristic methods such as the EM algorithm work well. This is because diverse or sparse cheaters are less correlated with the majority of the crowd and can be easily flagged as outliers using the classic methods. Despite this, our method demonstrates the most robust performance across a wide range of settings. Future work could explore integrating the principal’s samples of potential cheating signals into heuristic approaches like EM to enhance their performance and optimize detection.
@@
7 Conclusion
This work addresses the challenge of LLM-assisted cheating in crowdsourcing by developing a theoretically grounded quality measure for annotation tasks. Leveraging an extension of peer prediction, our method evaluates human information beyond what LLMs can generate, ensuring reliable human contributions. We establish theoretical guarantees under different cheating strategies and validate our approach empirically on real-world datasets. Our results demonstrate that our method effectively detects low-effort cheating while maintaining robustness across various conditions. Future work can focus on refining our method to better differentiate between harmful cheating and productive human-AI collaboration. This includes optimizing scoring mechanisms to recognize when LLMs are used as assistive tools rather than substitutes for human effort and developing adaptive detection techniques that adjust based on task characteristics and worker behavior.
Acknowledgment
This work was partly carried out at the DIMACS Center at Rutgers University. Jinlong Pang and Yang Liu are partially supported by the National Science Foundation (NSF) under grant IIS-2143895.
Appendix A Preliminary

### v2

@@ 5.2 Assumption Verification
The first two hypotheses correspond to 4.1 and 4.2 respectively and the other two hypotheses correspond to 4.5. More broadly, our results in this section provide valuable insights into the correlations between different LLMs and their correlations compared with human agents in the context of subjective labeling questions.
@@
For each considered LLM (as detailed in Section G.1), we prompt it to independently generate responses to each question in the dataset three times using the default temperature. These responses are treated as samples of the cheap signals $Z_{i}$, $Z_{j}$, and $Z$. Furthermore, we randomly select two human responses in the original dataset as samples of $X_{i}$ and $X_{j}$, respectively. Using these samples across all tasks, we estimate the conditioned joint distributions and then compute the conditional mutual information. We report the results for $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$ in Figure 2, while results for $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$ are deferred to Section G.3.
(a) Preference Alignment
(b) Hatefulness
(c) Toxicity
(d) Offensiveness
@@
Figure 2: The TVD mutual information between $X_{i}$ and $X_{j}$ (for “Human”, “Normal worker”, and “expert”) or $Z_{i}$ and $X_{j}$ (for remaining bars) conditioned on $Z$, i.e. $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$ or $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$. LLM names on the x-axis denote the models used to sample $Z$ while each bar in the same group represents the model used to sample $Z_{i}$. We use “Random Agents” as a reference, who randomly selects an answer to each question according to the prior.
@@
Hypotheses 1 and 2 generally do not hold.  In Figure 2, we observe that when $Z_{i}$ and $Z$ are sampled from different models, the mutual information between $Z_{i}$ and $X_{j}$ is often significantly higher than that expected from the random reporting (red bar). Even when $Z_{i}$ and $Z$ are independently sampled by the same model, conditioning on $Z$ reduces but does not eliminate this correlation. This indicates that even independently generated responses from the same LLM can remain dependent after conditioning on another independent sample, contradicting Hypothesis 1. As presented in Section G.3, this pattern continues to hold for $I_{\text{TVD}}(Z_{i};Z_{j}\mid Z)$, rejecting hypothesis 2.
@@
Hypothesis 3 and 4 generally hold when conditioning on the “right” model.  Encouragingly, the conditioned correlation between an LLM and a human agent is consistently weaker than that between two human agents when $Z_{i}$ and $Z$ come from the same model. For example, in Figure 2(a), the unconditional mutual information between GPT-4-generated labels and human labels exceeds that of two expert labels. However, after conditioning on GPT-4 responses, $I_{\text{TVD}}(X_{i};X_{j}\mid Z)$ becomes significantly higher than $I_{\text{TVD}}(Z_{i};X_{j}\mid Z)$. This pattern is consistent across all tested datasets, providing strong support for hypothesis 3. Results for Hypothesis 4 are deferred to the appendix.
@@
Our empirical results suggest that as of the time of this study and based on the subjective labeling tasks we tested, human agents exhibit distinct correlation patterns that no LLM replicates.
5.3 Detecting Low-effort Agents
@@
In this subsection, we focus on identifying low-effort agents with lazy-reporting strategies and comparing various methods based on their area under the ROC curve (AUC).
@@ 5.3.1 Simulating Noisy Crowds
Real-world crowdsourcing datasets do not include labels identifying whether an agent is exerting low effort or working diligently. To enable evaluation, we simulate low-effort agents and treat the original dataset labels as representing “full-effort” responses. Specifically, we replace a randomly selected fraction of agents in the original dataset with simulated low-effort agents adopting lazy-reporting strategies. To reflect the noisy nature of real-world crowdsourcing data, we simulate three types of low-effort agents, each corresponding to a common lazy-reporting strategy observed in practice yang2019modeling; becker2021crowd:
@@
    An $\alpha_{llm}$ fraction are LLM-reliant agents, who report the LLM-generated labels on all assigned questions.
@@
    An $\alpha_{r}$ fraction are random agents, who generate labels by independently sampling from the marginal distribution over labels for each assigned question.
@@
    An $\alpha_{b}$ fraction are biased agents, who report the dataset’s majority label on 90% of questions and choose uniformly at random on the rest.
Figure 3(a) shows the distribution of conditioned CA scores, where both the LLM-reliant agents’ responses and the principal’s signals are independently sampled from separate GPT-4 outputs. As shown, human workers achieve higher scores than any type of low-effort agent on average. However, some workers score below the average low-effort agent. One possible explanation is that a subset of workers in the dataset may also have exerted low effort. Due to these factors, no detection method can achieve perfect accuracy.
(a) Score distribution
(b) AUC
Figure 3: (a) Distribution of the conditioned CA scores with $(\alpha_{llm},\alpha_{r},\alpha_{b})=(0.1,0.05,0.05)$. (b) AUC scores of various methods as the fraction of LLM-reliant agents increases. In both panels, GPT-4 generates samples for $Z_{i}$, $Z_{j}$, and $Z$ using the preference alignment dataset.
@@ 5.3.2 Baseline Measurements
We introduce the key ideas of the following four baselines and defer the details to Section G.2.
@@
    We consider the Output Agreement (OA) mechanism that directly counts the frequency of agreements between two agents’ responses.
@@
    We use the Correlated Agreement (CA) mechanism as a baseline, which is introduced in Appendix B. The expected CA score is the unconditional mutual information $I_{\text{TVD}}(\hat{X}_{i};\hat{X}_{j})$.
@@
    Another common method to estimate the reliability of the noisy agents is the Expectation-Maximization (EM) algorithm under the DS model (dawid1979maximum). The algorithm iteratively updates a confusion matrix that captures the noise level of each agent. After convergence, we use the trace of each agent’s confusion matrix as a reliability score.
@@
    Lastly, we consider the Conditioned Agreement ($\text{OA}_{Z}$) which counts the frequency of agreements only when the agreement is different from the principal’s sample on that question.
We note that there are many more heuristic methods to estimate the reliability of crowdsourcing agents 4415269; embretson2013item; Piech2013. Many of these methods are variants of the EM algorithm under a different context-dependent model. We nonetheless only use the EM algorithm with the DS model to represent this line of methods because they all ignore the principal’s information $Z$ and thus are unlikely to perform well with LLM-corrupted datasets.
@@ 5.3.3 Results
Given a dataset $D$ and samples of the cheap signal $Z$, we use our method as well as the baselines to score each agent. We use the AUC score to measure the ability of each discussed metric to distinguish the simulated low-effort agents from the original human agents. In particular, let $I^{+}$ be the set of human agents and $I^{-}$ be the set of simulated low-effort agents. We compute the AUC score for each method as follows:
$\mathrm{AUC}=\frac{1}{|I^{+}|\cdot|I^{-}|}\sum_{i\in I^{+}}\sum_{j\in I^{-}}\left(\mathds{1}[S_{i}>S_{j}]+0.5\cdot\mathds{1}[S_{i}=S_{j}]\right).$
Figure 3(b) presents a comparison of different methods. For each value of $\alpha_{llm}$ ranging from 0 to $0.2$, we randomly sample $\alpha_{r}$ and $\alpha_{b}$ uniformly from $[0,0.2]$ and report the average and error bars over $50$ such trials. Our result suggests that classic methods, such as the EM algorithm and the CA mechanism, perform well when the number of LLM-reliant agents is low. However, their AUCs degrade significantly as the fraction of LLM-reliant agents increases. On the other hand, the performance of $\text{OA}_{Z}$, which scores agents based on the similarity between their reports and $Z$, tends to have large variance and performs poorly when there are few LLM-reliant agents. In contrast, our proposed conditioned CA mechanism has the most robust performance for the mixed crowds and has the lowest variance.
Table 1 summarizes the results from Figure 3(b), reporting the average AUC score and the bottom 10% quantile for each method on each dataset, where the average is w.r.t. every tested combination of $(\alpha_{llm},\alpha_{r},\alpha_{b})$. A high average AUC reflects strong overall performance, while a high 10% quantile indicates robustness to the worst-case scenarios. The results support our earlier claim that our method consistently demonstrates the highest robustness across all tested settings.
Table 1: Average AUC (and bottom 10% quantiles) across datasets and methods with GPT-4-reliant agents.
@@
We note that our experiments do not claim that our method outperforms all baselines in every scenario. Traditional methods like EM often perform better when the crowd consists entirely of random or biased agents. However, our method demonstrates the most robust performance across all settings due to its theoretical guarantees. This is particularly useful as, in practice, the designer usually has little clue about the composition of the crowds. In contrast, the baselines exhibit cases where the AUC drops below 0.5, meaning the algorithm completely misclassifies good workers as low-effort agents.
@@
First, Section G.4.1 presents analogous figures and tables for additional datasets and LLMs, further supporting the claims made in the main body. Second, we investigate how many questions the principal needs to sample for effective detection of low-effort agents (Section G.4.2). The results suggest that sampling LLM labels on 50% of the questions is typically sufficient to outperform the CA baseline. In cases where human and AI labels are well separated, sampling just 20% can suffice. Third, when the principal is uncertain about which LLMs the agents use, we propose a method that computes the CA score conditioned on each candidate LLM’s outputs and uses the minimum score across them as the final score (Section G.4.3). Finally, in Appendix F and Section G.5, we present a heuristic extension of our method to open-ended textual responses. This serves as a preliminary step toward generalizing our approach beyond labeling tasks, with further refinement and large-scale validation left for future work.
6 Conclusion
This work addresses the challenge of LLM contamination in crowdsourcing by developing a theoretically grounded quality measure for annotation tasks. Leveraging an extension of peer prediction, our method evaluates worker responses beyond what LLMs can generate, ensuring reliable human contributions. We establish theoretical guarantees under different low-effort strategies and validate our approach empirically on real-world datasets. Our results demonstrate that the proposed method effectively detects low-effort agents while maintaining robustness across various conditions. Future work can focus on refining our method to better differentiate between harmful LLM usage and productive human-AI collaboration. This includes optimizing scoring mechanisms to recognize when LLMs are used as assistive tools rather than substitutes for human effort and developing adaptive detection techniques that adjust based on task characteristics and worker behavior.
Broader Impact
On the positive side, our work can improve the robustness and fairness of data collection pipelines used in training AI systems, thereby enhancing the reliability of downstream decision-making or other usage of AI. However, there is a potential risk that insights into reporting strategies or detection weaknesses could be misused to evade quality control. We believe this risk is mitigated by focusing on defensive mechanisms and open discussion of vulnerabilities to promote transparency and better system design. Furthermore, like many crowdsourcing algorithms for label-quality control, our method operates without access to ground truth. Consequently, a low score does not always indicate a low-effort worker, while it may also reflect a diligent agent holding a minority opinion. Therefore, while such verification-free algorithms are highly scalable, they should be used with a human-in-the-loop, allowing the principal to carefully review low-scoring agents and prevent potential ethical concerns.
Acknowledgment
This work was partly carried out at the DIMACS Center at Rutgers University.
Appendix A Limitations
@@
We acknowledge several limitations of our method. First, our method is primarily designed for crowdsourcing datasets comprising subjective questions or tasks where human experts are considered the gold standard. If LLMs significantly surpass human performance, the correlations between human responses may no longer reflect the ground truth $Y$. This violates the assumption that, conditioned on $Y$, agents’ high-effort signals are independent. In such cases, applying our method could mistakenly reward incorrect correlations.
@@
Second, as discussed in our experiment section, our method does not outperform all baselines in every scenario. In particular, when LLM responses are diverse, e.g. when each agent relies on a different model or only a few agents use LLM assistance, heuristic methods such as the EM algorithm work well. This is because diverse or sparse low-effort agents are less correlated with the majority of the crowd and can be easily flagged as outliers using the classic methods. Despite this, our method demonstrates the most robust performance across a wide range of settings. Future work could incorporate the principal’s sampled LLM signals into heuristic methods such as EM to improve their performance and enhance detection under LLM contamination.
@@
Appendix B Preliminary
What do we know about the problem? Our work builds on advancements in information elicitation, particularly peer prediction. First introduced by miller2005eliciting, peer prediction mechanisms aim to incentivize truthful reporting without requiring ground truth verification. The core idea is to reward agents based on how well their responses predict those of others. The original mechanism relied on the principal’s knowledge of the correlation between agents’ signals (the information they observe). Later, dasgupta2013crowdsourced extended this idea, removing the need for such knowledge by assuming agents answer multiple i.i.d. binary questions. Subsequent works have expanded this framework to non-binary (but finite) signals [shnayder2016informed, kong2019information], infinite signals [schoenebeck2020learning], truthful guarantees with fewer questions [kong2020dominantly], heterogeneous agents [Agarwal2017], adversarial agents [rowdycrowds], and more general reporting strategies [zhangomni2023].